OpenStudy (anonymous):

pls help! Write an equation in point-slope form for the line parallel to -18x + 2y = 7 that contains (3, 1).

OpenStudy (anonymous):

solve for y 2y=18x+7 y=9x+7/2 m=9 m=-1/9

OpenStudy (anonymous):

thankyou :)

OpenStudy (anonymous):

y-3=-1/9(x-1) -9y-27=x-1 -9y-x=26

OpenStudy (anonymous):

you're welcome!

OpenStudy (anonymous):

wait sorry its parallel m is just 9 because i thought it said perpendicular sorry...

OpenStudy (anonymous):

oh..... i already submitted ur answer....its ok thx anyways

OpenStudy (anonymous):

this is for an online quiz? :O eek well i have a super headacher i shouldnt answer anymore question...sorry!!! omg...:(

OpenStudy (anonymous):

its ok :)